With the help of relevant examples, discuss any four costs and their implications for an organisation that embarks on a supply chain integration process.
Answer #1

There are various costs will be involved in the organization for supply chain process.

Below are the main four costs and their implication on organization which are involved in supply chain integration.

  1. Ordering or procurement cost:

Ordering/ Procurement cost are the cost’s that are involved in ordering or purchasing of any commodity or raw materials to produce the main or finished goods.

These are main cost to the any organization because most of the money is invested in purchasing the product.

Example: company XYZ purchase the 1000 units of mild steel, each unit costing $ 100 /-

Therefore the ordering cost = 1000*100 = $ 100000 /-

  1. Manufacturing cost:

The costs which are involved to produce the finish goods from the raw material is called as manufacturing cost.

Example: after purchasing the goods. Company XYZ manufacture good. The cost of manufacturing of each unit is $ 5

Therefore,

Manufacturing for 1000 units = 1000*5 = $ 5000 /-

  1. Inventory cost:

These are cost which are involved in keeping the manufactured goods within the organization, which include warehouse rent, electricity bills etc.

Example:

Company XYZ keeps those 1000 units of finished goods in the company warehouse. Ware house rent of $500 and electricity $ 150

Hence total inventory cost = 500+150 = $ 650

  1. Transportation cost:

These are the cost involved in transporting the finished goods till it reaching the customers. This involves fuel charge, road tolls or taxes.

Example: Company XYZ transport 1000 units of finished goods to customers by road, which cost $700/-

Implications:

All these above costs have direct implication on company profit and loss. Company can make profit by reducing the above cost as much as possible in a good and cheapest way. Increase in above cost might leads to lose the business.